The distance, x mm, moved by a particle is related to the time in seconds. We have

x = 1.2 sin(2t)

Find the distance travelled, the velocity and the acceleration of the particle at a time of t = 0.3 seconds.


1
Expert's answer
2021-06-07T14:55:30-0400
"x(t)=1.2\sin(2t)"

"v(t)=x'(t)=2.4\cos(2t)"

"a(t)=v'(t)=-4.8\sin(2t)"


"x(0.3)=1.2\sin(2(0.3))=1.2\sin(0.6) (mm)"

"\approx0.677571 (mm)"


"v(0.3)=2.4\cos(2(0.3))=2.4\cos(0.6) (mm/s)"

"\approx1.980805(mm/s)"


"a(0.3)=-4.8\sin(2(0.3))=-4.8\sin(0.6) (mm/s^2)"

"\approx-2.710284(mm/s^2)"